OpenStudy (ccastudentgrade10):

FAN AND MEDAL PLS HELP what is the equation in standard form of a parabola that models the values in the table? x = -2, 0, 4 f(x) =1, 5, -59 please xplain how you get it! this is a practice problem.

OpenStudy (welshfella):

y = ax^2 + bx + c is the general form of a parabola

OpenStudy (welshfella):

If you plug the given values of x and y into this you will get 3 equations in a,b and c which you can solve simultaneously

OpenStudy (ccastudentgrade10):

OHHHHHH i get it now thank you!!!!!

OpenStudy (welshfella):

for example x = 0 f(x) = 5 so we have a(0)^2 + 0b + c = 5 c = 5 so we have tha value of c already

OpenStudy (welshfella):

x = -2 , f(x) = 1:- 1 = a(-2)^2 - 2b + 5 4a - 2b = -4 is the second equation

OpenStudy (welshfella):

perform the same process with x = 4 and f(x) = -59 to get another equation in a and b Then you can solve the system of 2 equation to get the values of a and b